UT Nov 16 2021 - E/D Night split with PETS/PEPSI



Night Info:

  • Staring Config - MODS-Bino
  • Sunset 17:23
  • Twilight Ends 18:43
  • PEPSI/PETS starts UT 05hr (22hr local time) - reconfig for PEPSI at 21:45hr
  • PEPSI/PETS ends UT 10hr (03hr local time)
  • Reconfig for ?
  • Moon sets 04:24
  • Twilight Starts 05:26
  • Sunrise 06:46
  • ´╗┐MORNING: Prep LUCI-1 for mask exchange

Night Plan:

  • Run LUCI-software beta version w/updated RTD on OBS5
  • DONE MODS-Bino - M31 TOO - Wagner 1-1.5hr w/overheads
    • Any conditions
    • 18:30-18:55 - SpecStd Feige 110
    • 18:56-20:00 - M31 TOO
  • Reconfigure from MODS to LUCI - 20:00hr-20:15hr
  • LUCI-Bino Sub-frame readout testing 20:15hr-21:45hr
    • Any conditions
    • DONE Sub-frame readout software testing
    • DONE Test persistence (bright star outside of readout area0
    • Test ADI scripting
  • Reconfigure LUCI to PEPSI 21:45hr
  • DONE PEPSI/PETS 22hr-03hr
  • Reconfigure to MODS 03hr-03:15hr
  • DONE MODS Science 03:15-twilight
    • if more time needed for LUCI testing - continue in this block of time - need to prepare targets for later times
  • Poor Conditions BACKUP:
    • Local Mergers LUCI-Bino backup for 03:15hr - twilight


Staff:

  • OSA: D.G.Huerta
  • ISp: E. Solheid
  • ISA: B. Rothberg, R.T.Gatto
  • Mtn Manager: M. Wagner

Night Log

Closed Dome Activities

  • UT 22:58 (-1 day) - waking MODS and taking simSnaps
  • -MODS2 wake timed out
  • Retried and ran OK
  • simSnap ok
  • UT 23:12 (-1 day) - reconfiguring for MODS
  • UT 23:21 (-1 day) - starting up LUCI GUIs - L2 MOS came up unknown
    • initialize error on MOS unit
    • Verified retainer is locked, mask is clamped in FPU via switchbox
    • Strain gauges were on - zeroed out, and was able to initialize MOS
  • L1 0001-0002 2.51sec dark
  • L2 0001-0002 2.51sec dark
  • UT 23:32 (-1 day) - put MODS in observing mode and homed AGWs
  • Ready to go for observing
  • UT 00:14 - opening enclosure
  • UT 00:30 - some wisps of clouds overhead

RMWagner - M31 Nova TOO

  • UT 00:42 - pointing check
  • UT 00:45 - collimation check - can't find star yet
  • UT 01:07 - Collimation struggling a little - seeing jumping around, SX having the most trouble
  • UT 01:08 - sending preset to Feige 110
    • M1 offsetxy 0.295 11.374 rel
    • M2 offsetxy 4.786 8.645 rel
    • Tweak M2 offsetxy -0.869 -0.067 rel
    • Tweak M1 offsetxy -0.481 0.346 rel
  • UT 01:15 - starting spectra
  • guiders down to 0.7"-0.8" seing, some bumpts to 0.9"
    • UT 01:28 - sending preset to science field M31 Nova
    • Using 0.8" slit - as guiders were showing 0.7"-0.8" seing during SpecStd observations
    • Changed script to do 15sec thruslit first image, 60sec for remainder of acquisition
    • M1 offsetxy 0.151 10.053 rel
    • M2 offsetxy 4.629 7.875 rel
    • M1 looks OK
    • M2 looks OK
    • UT 01:40 - starting spectroscopy
    • seeing at start 0.71"-0.85"
    • UT 01:48 - seeing blew up a bit - up to 1.5" befor coming back down to 1"
    • UT 01:53 - back down to <1"
    • UT 02:09 - seeing in guiders down to 0.65"-0.75"
    • UT 02:18 - seeing dropping to 0.5"-0.6"
    • UT 02:27 - seeing still 0.5"-0.6"
    • UT 02:39 - seeing 0.56"-0.65"
    • UT 02:48 Reconfiguring to bino-LUCI

LUCI Sub-Frame Readout Testing

Sub-Frame test with cubes (not bright)
  • UT 03:01 - pointing and collimation check
  • UT 03:05 - swing and a miss - nothing in pointing check
  • UT 03:09 - pointing check
  • UT 03:12 - sending script for QSO
  • First Unusual (never before seen) error:
QSO Sub-frame readout test1918.02.15 20:45:28 u | error medium | obs5.mou | Luci ONE | RealTimeDisplay | ReadoutChangeListener.java#processJourna(87) | Empty file name was retrieved from journalizer into Real Time Display.
1918.02.15 20:45:28 u | error medium | obs5.mou | Luci ONE | RealTimeDisplay | ReadoutChangeListener.java#processJourna(87) | Empty file name was retrieved from journalizer into Real Time Display.

  • Acquisistion - step successful - field image, on-target image, can use to background subtract (or for bright targets might only need one image)
    • There will be two templates, one for pair subtraction and one for bright object (single exposure)
  • Setting up 512x2048 sub-frame
  • PROBLEMS/ISSUES RTD:
    • RTD still defaulting the first object in the stack
    • Getting error where click on the pixel map and spins - Jenny is NOT getting it on her OBS5 machine
    • Two sessions on same OBS5 can interfere with each other unless INTERLOP--->DISCONNECT FROM SAMP is selected (maybe this should be by default)
    • Saves a summed cube -->#0008 appears as compressed image in Aladding, but missing #0007
      • will save save an integrated file image along with CUBE into /Repository
      • eventually archiving will be turned off for integrated image
      • luci1.20211116.0007.fits 0.6264943737 100 0.626494
        luci1.20211116.0008.fits 0.6264943737 100 62.649437
        #0007 is the integrated cube, #0008 actual cube image - EXPTIME is correct for integrated image but it shows both DIT and NDIT
      • if one isn't saved to archive, would have gaps in the numbering scheme - this might be a good idea to avoid people thinking they have extra data
      • Will need to update fits header
    • Now PIXAMP works?! i clicked OUTSIDE of the image in the list of files and that seemed to fix it?
    • Race Conditions:
      • With two users on OBS5, Jenny has problems with #0028 LUCI-2 - it shows a quarter of an image, however, mine RTD shows a perfectly normal looking image
      • L2 #0028 - loads fine manually, Jenny testing it on her RTD - and it loaded OK
    • Getting a number of remote call errors in RTD - could it be conflicting sessions?
    • Additonal Errors in RTD
    • RTD_Error.jpg

Sub-frame with Bright Target Readout
  • Bright target Template has absorbs
  • In Track Mode
  • J.Power regenerating script for bright target sub-frame
    • May need to check backend
    • Code originated from longslit acquisition - has absorb commands in there - will need to be fixed in backend
  • Proceeding with script as is to test rest of template and test crosstalk from bright star outside the sub-frame reoudout
  • Sub-window in a box 256x256
  • UT 04:23 - seeing blew up to 1.5"-2.5"
  • 0,0 absolute both sides
  • Unsure if seeing crosstalk as seeing got so bad stars no longer too bright
    • will need to check files by eye tomorrow to look for crosstalk

Handover to PEPSI/PETS

  • UT 04:35 - switching to PETS/PEPSI
  • LUCIs in safe config
  • WASP-82 for 5 hours
  • UT 10:00 completed

MODS Science mid-z ULIRG

  • Configuring for bino-MODS
    • seeing has been <<1" and steady - moon setting soon - ideal for this target
    • clouds are approaching from the west but target will be rising in the east
  • UT 10:12 - homing AGWs (already in observing mode)
  • UT 10:16 - pointing check
  • UT 10:20 - collimation check
  • UT 10:23 - sending preset to G191B2B
    • M1 offsetxy -0.660 11.257 rel
    • M2 offsetxy 4.945 8.806 rel
    • M1 looks OK
    • M2 tweak offsetxy -1.110 -0.185 rel
  • UT 10:30 - starting spectra
  • UT 10:38 - M1R being slow on the readout
  • UT 10:43 - sending preset to FF0937
    • M1 offsetxy -0.433 10.109 rel
    • M2 offsetxy 3.693 7.943 rel
    • Tweak M1 -0.12" DETX
    • M1 is so much fainter than M2 in acquisition images -
  • UT 11:00 -starting MODS2 while M1 thru slit is taken
    • undoing tweak M1 +0.12" DETX
  • UT 11:03 - starting MODS1
  • UT 12:00 - seeing has been <1" - usually 0.7"-0.9"
  • UT 12;30 - seeing 0.7"
  • UT 12:38 - MODS-2 is done
  • UT 12:41 - MODS-1 is done
  • UT 12:43 - closing done - No cals.


-- BarryRothberg - 14 Nov 2021
Topic revision: r8 - 16 Nov 2021, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback